# and even simpler proofs of quantumness

### v1

@@ 1 Introduction
[…]
Is there a way to compile an arbitrary contextuality game into an operational “test of contextuality”?
Beyond demonstrating the presence of genuine quantum behaviour, non-local games can be employed to achieve a much more fine-grained control over the behaviour of untrusted quantum devices: for example, they allow a classical user to verify the correctness of full-fledged quantum computations, by interacting with two non-communicating quantum devices [RUV13, CGJV19]. Of course, any guarantee obtained via non-local games hinges on the physical (and non-falsifiable) assumption that the devices involved are non-communicating. To circumvent the need for this assumption, a lot of the attention in recent years has shifted to the computational setting. This exploration was kick-started by Mahadev’s seminal work [Mah18] showing, via cryptographic techniques, that the verification of quantum computations can be achieved with a single quantum device, under the assumption that the device is computationally bounded. She and her collaborators [BCM+21] later proposed what can be thought of as the analogue of a Bell/CHSH experiment with a single device—they proposed a simple test that an efficient quantum device can pass, but that an efficient classical device cannot. Since then, various works have proposed increasingly efficient “proofs of quantumness” in this setting [BKVV20, ACGH20, KMCVY22, KLVY23, AMMW22, BGKM+23a]. The goal of this line of work is to simplify these tests to the point that they can be implemented on current quantum devices. An experimental realisation of such a proof of quantumness would be a milestone for the field of quantum computation, as it would constitute the first efficiently verifiable demonstration of quantum advantage. Towards this goal, the second question that we consider in this work is:
@@ 1.1 Our results
[…]
@@
Any contextuality game can be compiled into a single prover, 2-round (i.e. 4-message) protocol such that the following two conditions hold (assuming standard cryptographic assumptions):
@@
We require that no PPT prover can distinguish between the two cases with non-negligible advantage. We show that an oblivious pad can be realised based on ideas from [BCM+21] in the quantum random oracle model. [^12] We do not foresee any obstacle in obtaining a construction in the plain model, at the cost of constantly many additional rounds of interaction. Crucially, while in KLVY the second phase of the protocol happens on a different subsystem, the oblivious pad is what allows us to carry out the second phase on the same subsystem, as depicted in Figure 2.
@@
Our second contribution streamlines the ideas introduced to prove LABEL:Thm:infMain1 in order to obtain a 2-round “proof of quantumness” relying on the classical hardness of the Learning-with-Errors (LWE) problem [Reg09]. Our construction has the main advantage of being simpler than existing ones in the literature, in the sense explained below. Our proof of quantumness makes use of Noisy Trapdoor Claw-Free functions (NTCFs), introduced in [BCM+21] (but it does not require the NTCFs to have an “adaptive hardcore bit” property). It relies on the particular structure of the “encrypted CNOT operation” introduced in Mahadev’s QFHE scheme [Mah20]. We informally state our result, but we defer the construction to the technical overview.
@@
To the best of our knowledge, the only aspect in which our proof of quantumness compares unfavourably with existing ones, e.g. [KMCVY22], is that, based on current knowledge of constructions of NTCFs, our proof of quantumness requires a construction from LWE (in order to implement the “encrypted CNOT” procedure), whereas [KMCVY22] has the flexibility that it can be instantiated using any TCF, e.g. based on Diffie-Hellman or Rabin’s function. While the latter are more efficient to implement, they also generally require a larger security parameter (inverting Rabin’s function is as hard as factoring, whereas breaking LWE is as hard as worst-case lattice problems). Hence, it is currently still unclear which route is closer to a realisation at scale.
@@ 1.2 Future Directions
  • Better contextuality compilers. Focusing on compilers for contextuality, the following important aspect remains to be strengthened. The current compiler does not in general achieve quantum soundness (in the sense that there are contextuality games in which a QPT prover can do much better than the “compiled” quantum value from LABEL:Thm:infMain1). Interestingly, recent works show that KLVY does satisfy quantum soundness for certain families of non-local games [NZ23, BGKM+23b, CMM+24].
  • Oblivious Pauli pad. We think that the new functionality of the oblivious Pauli pad has the potential to be useful elsewhere, and we leave this exploration to future work. A related question is to realise the oblivious Pauli pad in the plain model (without increasing the number of rounds).
  • More efficient “encrypted CNOT”. Turning to proofs of quantumness, the broad goal is of course to simplify these even further towards experimental implementations. One concrete direction in which our proof of quantumness could be simplified is the following. Currently, we use an NTCF that supports the “encrypted CNOT” operation from [Mah20]. However, our only requirement is that the NTCF satisfies the potentially weaker property that it hides a bit in the xor of the first bit of a pair of preimages (see LABEL:Def:ntcf-2). This is because we only need to perform the “encrypted CNOT” operation once, and not repeatedly as part of a full-fledged homomorphic computation. It would be interesting to see if the weaker requirement could be achieved by simpler claw-free functions (from LWE or other assumptions, like DDH or factoring).
@@ Acknowledgements
We thank Ulysse Chabaud, Mauro Morales and Thomas Vidick for their feedback on early drafts of this work. We thank Yusuf Alnawakhtha, Alexandru Gheorghiu, Manasi Mangesh Shingane, and Urmila Mahadev for various helpful discussions. ASA acknowledges support from the U.S. Department of Defense through a QuICS Hartree Fellowship, IQIM, an NSF Physics Frontier Center (GBMF-1250002) and MURI grant FA9550-18-1-0161. Part of the work was carried out while ASA was visiting the Simons Institute for the Theory of Computing. AC acknowledges support from the UK National Quantum Computing Centre through the Quantum Advantage Pathfinder (QAP) research programme (EP/X026167/1).
@@ 2.2 Contextuality
(a) Memoryless Device
(b) A memoryless operational test, illustrated using the Peres-Mermin magic square.
Figure 5: The folklore memoryless interpretation of contextuality.
@@
Our construction follows the same general approach of enforcing separation in “time” instead of in “space”, but is a radical departure from the attempt described above. We only assume that the device is computationally bounded, and we use cryptographic techniques to construct an operational test, which we describe in the next sections.
@@ 2.3 Quantum Fully Homomorphic Encryption (QFHE)
    $\mathsf{Gen}$ is a classical algorithm, while $\mathsf{Enc},\mathsf{Dec}$ become classical algorithms when their inputs are classical. In particular, this means that classical inputs are encrypted into classical ciphertexts. This property is essential when the scheme is deployed in protocols involving classical parties, as will be the case here. [^15] The first schemes to satisfy this property appeared recently in the breakthrough works [Mah20, Bra18].
@@
    Encryption of a state $\left|\psi\right\rangle$ takes the form $(X^{x}Z^{z}\left|\psi\right\rangle,\widehat{xz})\leftarrow\mathsf{Enc}_{\mathsf{sk}}(\left|\psi\right\rangle)$, where $X^{x}$ applies a Pauli $X$ to the $i$-th qubit based on the value of $x_{i}$, and $Z^{z}$ is defined similarly. $\widehat{xz}$ is an encryption of the pad $xz$.
All known constructions of QFHE schemes satisfying property 1 also satisfy properties 2 and 3. The KLVY compiler, as one might guess, relies on property 2. Our compiler relies on property 3.
@@ 2.4 The KLVY Compiler
  • Round 2. Send an encryption $c_{y}$ of “Bob’s question”, and get an encrypted answer $c_{b}$ in the clear. Decrypt the answers using $\mathsf{sk}$, and accept if $\mathsf{pred}(a,b,x,y)=1$.
The honest prover prepares the entangled state |$\ket{\psi}_{AB}$ corresponding to the optimal quantum strategy for the non-local game. It uses $\mathsf{QFHE}$’s $\mathsf{Eval}$ algorithm on subsystem $A$ to answer question $x$ according to Alice’s optimal strategy, and answers question $y$ in the clear using Bob’s optimal strategy on subsystem $B$. More formally, they proceed as follows.
@@ 2.5 Contribution 1 — A Computational Test of Contextuality
The honest prover’s state after Round 1 is encrypted and has the form $(X^{x}Z^{z}\ket{\psi_{(q_{1},a_{1})}},\widehat{xz})$ where $\widehat{xz}$ denotes a classical encryption of the strings $xz$ (using Property 3 of the $\mathsf{QFHE}$ scheme), as detailed below.
@@
The oblivious $\mathbf{U}$ pad is a tuple of algorithms $(\mathsf{Gen},\mathsf{Enc},\mathsf{Dec})$ where $\mathsf{Gen}$ and $\mathsf{Dec}$ are PPT. Let $(\mathsf{pk},\mathsf{sk})\leftarrow\mathsf{Gen}(1^{\lambda})$ be the public and the secret keys generated by $\mathsf{Gen}$. Encryption takes the form $(U_{k}\left|\psi\right\rangle,s)\leftarrow\mathsf{Enc}_{\mathsf{pk}}(\left|\psi\right\rangle)$, where $k=\mathsf{Dec}_{\mathsf{sk}}(s)$. Notice the similarity with the post-measurement state in the discussion above (we will return to this in a moment). The security requirement is that no PPT algorithm can win the following security game with probability non-negligibly greater than $1/2$.
@@
We expect that, by allowing a larger constant number of rounds of interaction, it should be possible to realise the oblivious pad functionality in the plain model.
@@
Below, we summarise our compiler. Since we now have ${\mathsf{Gen},\mathsf{Enc},\mathsf{Dec}}$ algorithms for both $\mathsf{OPad}$ and $\mathsf{QFHE}$, we use prefixes such as $\mathsf{OPad}.\mathsf{Enc}$ to refer to $\mathsf{Enc}$ associated with $\mathsf{OPad}$ to avoid confusion.
@@
Consider a contextuality game with contexts of size $2$. Let ${\rm valNC}$ and ${\rm valQu}$ be its non-contextual and quantum values respectively. The compiled game (as described above) satisfies the following:
@@
Proof sketch: Suppose ${\cal A}$ is a PPT algorithm that wins with probability non-negligibly greater than $\frac{1}{2}(1+{\rm valNC})$. Observe that one can associate a “deterministic assignment” corresponding to ${\cal A}$, conditioned on some fixed first round messages, as follows: simply rewind ${\cal A}$ to learn answers to all possible second round questions, obtaining an assignment $\tau:Q\to A$, mapping questions to answers. Let us write $\tau_{q_{1}}$ to make the dependence of the assignment on the first question more explicit (note that the assignment depends on the encrypted question $c_{q_{1}}$ as well as the encrypted answer $c_{a_{1}}$). For the purpose of this overview, suppose also that ${\cal A}$ is consistent, i.e. if $q_{1}=q_{2}$, then $a_{1}=a_{2}$ (note that this in particular ensures that, when $q_{1}=q_{2}$, ${\cal A}$ wins with probability 1. One can show that an adversary that is not consistent can be turned into an adversary that is consistent and wins with at least the same probability). Now, to win with probability more than $\frac{1}{2}(1+{\rm valNC})$, it must be that the $\tau_{q_{1}}$’s are different for different $q_{1}$’s. Otherwise, ${\cal A}$’s strategy is just a convex combination of deterministic assignments and this by definition cannot do better than ${\rm valNC}$ when $q_{1}\neq q_{2}$. But if the distribution over $\tau_{q_{1}}$s and $\tau_{q_{1}^{\prime}}$s is different for at least some $q_{1}\neq q_{1}^{\prime}$, then one is able to distinguish $\mathsf{QFHE}$ encryptions of $q_{1}$ from those of $q_{1}^{\prime}$. Thus, as long as the $\mathsf{QFHE}$ scheme is secure, no PPT algorithm can win with probability non-negligibly greater than $\frac{1}{2}(1+{\rm valNC})$.
@@
Consider an arbitrary contextuality game, and let ${\rm valNC}$ and ${\rm valQu}$ be its non-contextual and quantum values respectively. The compiled game, obtained via the $(|C|-1,1)$ compiler, satisfies the following:
@@
The compiled game, obtained via the $(|C|,1)$ compiler, satisfies the following:

### v2

@@ 1 Introduction
Bell non-locality may be viewed as a special case of the more general phenomenon of quantum contextuality [Spe60, KS67], which says that such correlations can occur even when the measurements are not necessarily on separate quantum systems, but are merely “compatible” (i.e. commuting). Contextuality has a long tradition in the foundations of quantum mechanics [BCG+22]. However, unlike a non-local game, a more general contextuality game [^5] We choose to use the term contextuality “game” here to preserve the analogy with a non-local game. However, in the contextuality literature, the more commonly used term is contextuality “scenario”. We refer the reader to the technical overview (LABEL:Subsec:tech-overview-contextuality) for details. does not in general have a corresponding operational test. By an operational test, we mean a test that can be carried out on a device by interacting with it classically, and importantly, without having to make bespoke assumptions about its inner workings. Thus, even though some contextuality games are even simpler than non-local games [KCBS08], a satisfactory approach to compiling arbitrary contextuality games into operational “tests of contextuality” is missing. This is not for lack of trying—numerous attempts have been made that inevitably have to either resort to strong assumptions about the quantum hardware or assumptions that are hard to enforce in practice, such as the device being essentially memoryless. [^6] These assumptions are typically referred to as “loopholes” in the literature: [LLS+11, UZZ+13, JRO+16, ZKK+17, MZL+18, LMZ+18, ZXX+19, UZZ+20, WZL+22, HXA+23, LMX+23]. Thus, one of the central open questions in the foundations of quantum mechanics is:
Is there a way to compile an arbitrary contextuality game into an “operational test of contextuality”?
Beyond demonstrating the presence of genuine quantum behaviour, non-local games can be employed to achieve a much more fine-grained control over the behaviour of untrusted quantum devices: for example, they allow a classical user to verify the correctness of full-fledged quantum computations, by interacting with two non-communicating quantum devices [RUV13, CGJV19]. Of course, any guarantee obtained via non-local games hinges on the physical (and non-falsifiable) assumption that the devices involved are non-communicating. To circumvent the need for this assumption, a lot of the attention in recent years has shifted to the computational setting. This exploration was kick-started by Mahadev’s seminal work [Mah18] showing, via cryptographic techniques, that the verification of quantum computations can be achieved with a single quantum device, under the assumption that the device is computationally bounded. She and her collaborators [BCM+21] later proposed what can be thought of as the analogue of a Bell/CHSH experiment with a single device—they proposed a simple test that an efficient quantum device can pass, but that an efficient classical device cannot. Since then, various works have proposed increasingly efficient “proofs of quantumness” in this setting [BKVV20, ACGH20, KMCVY22, KLVY23, AMMW22, BGKM+23]. The goal of this line of work is to simplify these tests to the point that they can be implemented on current quantum devices. An experimental realisation of such a proof of quantumness would be a milestone for the field of quantum computation, as it would constitute the first efficiently verifiable demonstration of quantum advantage. Towards this goal, the second question that we consider in this work is:
@@ 1.1 Our results
Our first contribution is a positive answer to the first question. We show that, using cryptographic techniques, an arbitrary contextuality game can be compiled into an “operational test of contextuality” involving a single quantum device, where the only assumption is that the device is computationally bounded. [^7] In the contextuality jargon, one might say that all “loopholes” are being replaced by a computational assumption. A contextuality game involves a single player and a referee (unlike non-local games that always involve more than one player). In an execution of the game, the referee asks the player to measure a context—a set of commuting observables—and the player wins if the measurement outcomes satisfy certain constraints. Importantly, a given observable may appear in multiple contexts. If the player uses a strategy where the values of these observables are “predetermined” (which is the analogue of a “local hidden variable” strategy in the non-local game setting), then the highest winning probability achievable is referred to as the non-contextual value of the game, denoted by ${\rm valNC}$ [^8] We emphasise that a non-contextual strategy is such that, if an “observable” appears in multiple contexts, then the same predetermined value should be returned by the player for that observable in all contexts in which it appears. This is precisely the difficulty in realising an “operational test of contextuality”: how does the referee enforce that the player is consistent across contexts? On the other hand, if the player uses a quantum strategy, then the highest winning probability achievable is the quantum value, denoted by ${\rm valQu}$ (which is strictly greater than ${\rm valNC}$ for contextuality games of interest). [^9] A quantum strategy is such that, if an observable appears in multiple contexts, then the same observable should be measured to obtain the answer in all such contexts. We formalise what we mean by an “operational test of contextuality” by introducing a property called faithfulness. Intuitively, this requires that the test has a strong correspondence to some contextuality game. We defer more precise definitions to LABEL:Subsec:criteria. We show the following.
@@
Any contextuality game $\mathsf{G}$ can be compiled into a single prover, 2-round (i.e. 4-message) operational test of contextuality, under standard cryptographic assumptions. More precisely, the test is faithful to $\mathsf{G}$ and satisfies the following:
@@
We require that no PPT prover can distinguish between the two cases with non-negligible advantage. [^12] Note that there is a QPT prover that can distinguish these two cases with non-negligible advantage. We show that an oblivious pad can be realised based on ideas from [BCM+21] in the quantum random oracle model. Crucially, while in KLVY the second phase of the protocol happens on a different subsystem, the oblivious pad is what allows us to carry out the second phase on the same subsystem, as depicted in Figure 2.
@@
Our second contribution streamlines the ideas introduced to prove LABEL:Thm:infMain1 in order to obtain a 2-round proof of quantumness relying on the classical hardness of the Learning-with-Errors (LWE) problem [Reg09]. Our construction has the main advantage of being simpler than existing ones in the literature, in the sense explained below. Our proof of quantumness makes use of Noisy Trapdoor Claw-Free functions (NTCFs), introduced in [BCM+21] (but it does not require the NTCFs to have an “adaptive hardcore bit” property). It relies on the particular structure of the “encrypted CNOT operation” introduced in Mahadev’s QFHE scheme [Mah20]. We informally state our result, but we defer the construction to the technical overview.
@@
To the best of our knowledge, the only aspect in which our proof of quantumness compares unfavourably with existing ones, e.g. [KMCVY22], is that, based on current knowledge of constructions of NTCFs, our proof of quantumness requires a construction from LWE (in order to implement the “encrypted CNOT” procedure), whereas [KMCVY22] has the flexibility that it can be instantiated using any TCF, e.g. based on Diffie-Hellman or Rabin’s function. While the latter are more efficient to implement, they also generally require a larger security parameter (inverting Rabin’s function is as hard as factoring, whereas breaking LWE is as hard as worst-case lattice problems). Hence, it is currently still unclear which route is closer to a realisation at scale. In particular, we note that instantiating the construction of [KMCVY22] with the $x^{2}\bmod N$ function by relying on a novel multiplication algorithm leads to a very efficient quantum circuit for the prover as shown in [KMY24]. It is plausible that future improvements to our proof of quantumness might yield constructions from a broader class of hardness assumptions than LWE (e.g. Ring-LWE), which would likely yield a further improvement in concrete efficiency [^13] An upcoming work by one of the authors in fact improves and instantiates our template using Ring-LWE..
@@ 1.2 Future Directions
  • Better contextuality compilers. Focusing on compilers for contextuality, the following important aspect remains to be strengthened. The current compiler does not in general achieve quantum soundness (in the sense that there are contextuality games in which a QPT prover can do much better than the “compiled” quantum value from LABEL:Thm:infMain1). Interestingly, recent works show that KLVY does satisfy quantum soundness for certain families of non-local games [NZ23, BGKM+23, CMM+24].
[…]
  • More efficient “encrypted CNOT”. Turning to proofs of quantumness, the broad goal is of course to simplify these even further towards experimental implementations. One concrete direction in which our proof of quantumness could be simplified is the following. Currently, we use an NTCF that supports the “encrypted CNOT” operation from [Mah20]. However, our only requirement is that the NTCF satisfies the potentially weaker property that it hides a bit in the xor of the first bit of a pair of preimages. This is because we only need to perform the “encrypted CNOT” operation once, and not repeatedly as part of a full-fledged homomorphic computation. It would be interesting to see if the weaker requirement could be achieved by simpler claw-free functions (from LWE or other assumptions, like DDH or factoring).
@@ Acknowledgements
We are thankful to an anonymous QCrypt reviewer for their observations about the relation between the oblivious pad and non-interactive proofs of quantumness, and for pointing out the limitations of not formalising what it means for a test to be an operational test of contextuality.
We thank Ulysse Chabaud, Mauro Morales and Thomas Vidick for their feedback on early drafts of this work. We thank Yusuf Alnawakhtha, Alexandru Gheorghiu, Manasi Mangesh Shingane, and Urmila Mahadev for various helpful discussions. ASA acknowledges support from the U.S. Department of Defense through a QuICS Hartree Fellowship, IQIM, an NSF Physics Frontier Center (GBMF-1250002) and MURI grant FA9550-18-1-0161. Part of the work was carried out while ASA was visiting the Simons Institute for the Theory of Computing. AC acknowledges support from the National Science Foundation grant CCF-1813814, from the AFOSR under Award Number FA9550-20-1-0108 and from the Engineering and Physical Sciences Research Council through the Hub in Quantum Computing and Simulation grant (EP/T001062/1) and the Quantum Advantage Pathfinder (QAP) research programme (EP/X026167/1). KB is supported by A*STAR C230917003.
@@ 2.2 Contextuality
Figure 5: The folklore memoryless interpretation of contextuality: a memoryless device (left) and a memoryless operational test, illustrated using the Peres-Mermin magic square (right).
@@
Our construction follows the same general approach of enforcing separation in “time” instead of in “space”, but is a radical departure from the attempt described above. In the subsequent sections, we only assume that the device is computationally bounded, and use cryptographic techniques to construct an “operational test of contextuality”.
Remark 3 (Criteria for being an Operational Test of Contextuality).
We have not yet formally defined what we mean by an “operational test of contextuality”. Consider any test that serves as a proof of quantumness. Intuitively, we require that this test, additionally, is faithtful to some contextuality game $\mathsf{G}$, i.e. it satisfies the following two properties:
  • The test involves asking the prover questions corresponding to $\mathsf{G}$, possibly under some “encoding”. The test is, however, allowed to involve other messages unrelated to $\mathsf{G}$.
  • Whether the test passes or fails is determined solely by evaluating the predicate corresponding to $\mathsf{G}$, using the “decoded” questions and answers.
Formalising these requirements is a bit more involved and is deferred to LABEL:Subsec:criteria. However, this intuitive notion suffices for now, as will be clear from our construction in the subsequent sections.
@@ 2.3 Quantum Fully Homomorphic Encryption (QFHE)
    $\mathsf{Gen}$ is a classical algorithm, while $\mathsf{Enc},\mathsf{Dec}$ become classical algorithms when their inputs are classical. In particular, this means that classical inputs are encrypted into classical ciphertexts. This property is essential when the scheme is deployed in protocols involving classical parties, as will be the case here. [^16] The first schemes to satisfy this property appeared in the breakthrough works [Mah20, Bra18].
@@
    Encryption of a state $\left|\psi\right\rangle$ takes the form $(X^{x}Z^{z}\left|\psi\right\rangle,\widehat{xz})\leftarrow\mathsf{Enc}_{\mathsf{sk}}(\left|\psi\right\rangle)$, where $X^{x}$ applies a Pauli $X$ to the $i$-th qubit based on the value of $x_{i}$, and $Z^{z}$ is defined similarly. $\widehat{xz}$ is a classical $\mathsf{FHE}$ encryption of the pad $xz$.
All known constructions of QFHE schemes satisfying property 1 also satisfy properties 2 and 3. The KLVY compiler, as one might guess, relies on property 2. Our compiler relies on property 3.171717 Strictly speaking, we require the $\mathsf{FHE}$ scheme in property 3 to be public-key. This is used to establish the faithfulness condition (detailed in LABEL:Def:OperationalTestOfContextuality and LABEL:Subsec:FHE-Scheme). We neglect this here for readability.
@@ 2.4 The KLVY Compiler
  • Round 2. Send “Bob’s question” $y$, and get his answer $b$ in the clear. Decrypt Alice’s response using $\mathsf{sk}$, and accept if $\mathsf{pred}(a,b,x,y)=1$.
The honest prover prepares the entangled state |$\ket{\psi}_{AB}$ corresponding to the optimal quantum strategy for the non-local game. It uses $\mathsf{QFHE}$’s $\mathsf{Eval}$ algorithm on subsystem $A$ to answer question $x$ according to Alice’s optimal strategy, and answers question $y$ in the clear using Bob’s optimal strategy on subsystem $B$. More formally, they proceed as in LABEL:Fig:KLVY.
@@
Figure 6: The KLVY compiler [KLVY23] is illustrated above: it takes any two-player non-local game and converts it into a proof of quantumness.
@@ 2.5 Contribution 1 — A Computational Test of Contextuality
The honest prover’s state after Round 1 is encrypted and has the form $(X^{x}Z^{z}\ket{\psi_{(q_{1},a_{1})}},\widehat{xz})$ where $\widehat{xz}$ denotes a classical encryption of the strings $xz$ (using Property 3 of the $\mathsf{QFHE}$ scheme), as detailed below in Figure 7.
@@
Figure 7: The figure illustrates that, a priori, it is unclear how to generalise KLVY to compile contextuality games.
@@
[…]
@@
    Repeat steps (i) and (ii) on |$\ket{\psi_{Z}}$, but in the Hadamard basis! This results in strings $y^{\prime}$ and $d^{\prime}$, as well as a leftover qubit state |$\ket{\psi_{XZ}}=$
    $X^{d^{\prime}\cdot(x_{0}^{\prime}\oplus x_{1}^{\prime})+H(x_{0}^{\prime})+H(x_{1}^{\prime})}Z^{d\cdot(x_{0}\oplus x_{1})+H(x_{0})+H(x_{1})}\ket{\psi}\,,$
@@
We summarise our compiler in Figure 9. Since we now have ${\mathsf{Gen},\mathsf{Enc},\mathsf{Dec}}$ algorithms for both $\mathsf{OPad}$ and $\mathsf{QFHE}$, we use prefixes such as $\mathsf{OPad}.\mathsf{Enc}$ to refer to $\mathsf{Enc}$ associated with $\mathsf{OPad}$ to avoid confusion.
@@
Figure 9: Our contextuality game compiler. It takes a contextuality game given by $(Q,A,C^{{\rm all}},\mathsf{pred},{\cal D}_{\rm contexts})$, a $\mathsf{QFHE}$ scheme and an $\mathsf{OPad}$ and returns an operational test of contextuality.
@@
Consider a contextuality game $\mathsf{G}$ with contexts of size $2$. Let ${\rm valNC}$ and ${\rm valQu}$ be its non-contextual and quantum values respectively. The compiled game (as described above) is faithful to $\mathsf{G}$. In particular, it satisfies the following:
@@
Proof sketch: The faithfulness condition as discussed in LABEL:Rem:criteria is evidently satisfied by the compiled game. [^19] The formal notion, as stated later in LABEL:Subsec:criteria, is also satisfied but the details are deferred. Suppose ${\cal A}$ is a PPT algorithm that wins with probability non-negligibly greater than $\frac{1}{2}(1+{\rm valNC})$. Observe that one can associate a “deterministic assignment” corresponding to ${\cal A}$, conditioned on some fixed first round messages, as follows: simply rewind ${\cal A}$ to learn answers to all possible second round questions, obtaining an assignment $\tau:Q\to A$, mapping questions to answers. Let us write $\tau_{q_{1}}$ to make the dependence of the assignment on the first question more explicit (note that the assignment depends on the encrypted question $c_{q_{1}}$ as well as the encrypted answer $c_{a_{1}}$). For the purpose of this overview, suppose also that ${\cal A}$ is consistent, i.e. if $q_{1}=q_{2}$, then $a_{1}=a_{2}$ (note that this in particular ensures that, when $q_{1}=q_{2}$, ${\cal A}$ wins with probability 1. One can show that an adversary that is not consistent can be turned into an adversary that is consistent and wins with at least the same probability). Now, to win with probability more than $\frac{1}{2}(1+{\rm valNC})$, it must be that the $\tau_{q_{1}}$’s are different for different $q_{1}$’s. Otherwise, ${\cal A}$’s strategy is just a convex combination of deterministic assignments and this by definition cannot do better than ${\rm valNC}$ when $q_{1}\neq q_{2}$. But if the distribution over $\tau_{q_{1}}$s and $\tau_{q_{1}^{\prime}}$s is different for at least some $q_{1}\neq q_{1}^{\prime}$, then one is able to distinguish $\mathsf{QFHE}$ encryptions of $q_{1}$ from those of $q_{1}^{\prime}$. Thus, as long as the $\mathsf{QFHE}$ scheme is secure, no PPT algorithm can win with probability non-negligibly greater than $\frac{1}{2}(1+{\rm valNC})$.
@@
Consider an arbitrary contextuality game $\mathsf{G}$, and let ${\rm valNC}$ and ${\rm valQu}$ be its non-contextual and quantum values respectively. The compiled game, obtained via the $(|C|-1,1)$ compiler, is faithful to $\mathsf{G}$ and satisfies the following:
@@
The compiled game, obtained via the $(|C|,1)$ compiler, is also faithful to $\mathsf{G}$, and satisfies the following: